# Algorithmic contiguity from low-degree conjecture and applications in correlated random graphs

### v1

@@ 4 Detection in correlated SBMs
In this section, we will use the framework we established in Section 2 to show the hardness of detection between correlated SBMs and independent SBMs. The main results of this section is incorporated as follows.
@@
Assuming Conjecture 2.2, when $\epsilon^{2}\lambda s<1-\delta$, $s<\sqrt{\alpha}-\delta$ and $\lambda>\lambda_{0}$, there is no algorithms with polynomial running time that can strongly distinguish $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and two independent $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$.
@@
Thus, this algorithm $\mathcal{A}$ strongly distinguish $\mu_{\mathsf{CorSBM}}$ and $\mu_{\mathsf{IndER}}$. This contradicts Conjecture 2.2 and the low-degree hardness established in [9, Theorem 1.3]. ∎
The rest part of this section is devoted to the proof of Theorem 4.1. For notational simplicity, in the following we will only prove the case where $K=1$ and the proof for general $K$ is similar. Note that $\epsilon^{2}\lambda<1$ and $\lambda>\lambda_{0}$ implies that $\epsilon<\epsilon_{0}=\lambda_{0}^{-1/2}$. We choose $\lambda_{0}$ to be a sufficient large constant such that
${}(\epsilon^{2}\lambda)\cdot\tfrac{(k-1)\sqrt{1-\epsilon}+\sqrt{1+\epsilon(k-1)}}{k}\geq 1-\delta/2\mbox{ for all }\epsilon<\epsilon_{0}=\lambda_{0}^{-1/2}\,.$ (4.2)
@@
$\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O(1)\mbox{ and }\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\big{)}=O(1)\,.$
Indeed, it has been shown in [32] that $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O(1)$ provided with (4.3). It remains to show that under (4.3) we have (note that now $\mathbb{P}$ is the planted measure)
${}\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\big{)}=\sup_{f\in\mathcal{P}_{D}}\frac{\mathbb{E}_{\mathbb{Q}}[f]}{\sqrt{\mathbb{E}_{\mathbb{P}}[f^{2}]}}=O_{\delta}(1)\,.$ (4.4)
We point out that our approach to proving (4.4) is based on the the work [52]. While the authors of [52] mainly focus on showing the low-degree hardness for community recovery problems, extending their approach to bound the left hand side of (4.4) is relatively standard. To this end, define
@@
It is well known in [32] that $\{\phi_{S}:S\Subset\mathcal{K}_{n},|E(S)|\leq D\}$ constitutes a standard orthogonal basis of $\mathcal{P}_{D}$ under $\mathbb{Q}$. Thus, each $f\in\mathcal{P}_{D}$ can be written as
@@
Regarding (4.18), it suffices to show that there exists $Mu^{\top}=c$ and $\|u\|=O(1)$. Note that $Mu^{\top}=c$ is equivalent to
@@
To prove (4.4), it suffices to show the following lemma

### v2

@@ 4 Detection in correlated SBMs
In this section, we will use the framework we established in Section 2 in stochastic block models below KS-threshold. The main results of this section is incorporated as follows.
@@
Assuming Conjecture 2.2, when $\epsilon^{2}\lambda s<1-\delta$, $s<\sqrt{\alpha}-\delta$ and $\lambda>\lambda_{0}(\delta,k)$ there is no algorithms with polynomial running time that can strongly distinguish $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and two independent $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$.
@@
Thus, this algorithm $\mathcal{A}$ strongly distinguish $\mu_{\mathsf{CorSBM}}$ and $\mu_{\mathsf{IndER}}$. This contradicts Conjecture 2.2 and the low-degree hardness established in [CDGL24+, Theorem 1.3]. ∎
The rest part of this section is devoted to the proof of Theorem 4.1. For notational simplicity, in the following we will only prove the case where $K=1$ and the proof for general $K$ is similar. Note that $\epsilon^{2}\lambda<1-\delta$ and $\lambda>\lambda_{0}$ implies that $\epsilon<\epsilon_{0}=\lambda_{0}^{-1/2}$. We choose $\lambda_{0}=\lambda_{0}(\delta,k)$ to be a sufficient large constant such that
${}(1-\delta)^{-1}\cdot\tfrac{(k-1)\sqrt{1-\epsilon}+\sqrt{1+\epsilon(k-1)}}{k}\geq(1-\delta/2)^{-1}\mbox{ for all }\epsilon<\epsilon_{0}=\lambda_{0}^{-1/2}\,.$ (4.2)
@@
$\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta,k}(1)\mbox{ and }\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\big{)}=O_{\delta,k}(1)\,.$
Indeed, it has been shown in [HS17] that $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta,k}(1)$ provided with (4.3). It remains to show that under (4.3) we have (note that now $\mathbb{P}$ is the planted measure)
${}\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\big{)}=\sup_{f\in\mathcal{P}_{D}}\frac{\mathbb{E}_{\mathbb{Q}}[f]}{\sqrt{\mathbb{E}_{\mathbb{P}}[f^{2}]}}=O_{\delta,k}(1)\,.$ (4.4)
We point out that our approach to proving (4.4) is based on the the work [SW25]. To this end, define
@@
It is well known in [HS17] that $\{\phi_{S}:S\Subset\mathcal{K}_{n},|E(S)|\leq D\}$ constitutes a standard orthogonal basis of $\mathcal{P}_{D}$ under $\mathbb{Q}$. Thus, each $f\in\mathcal{P}_{D}$ can be written as
@@
Regarding (4.18), it suffices to show that there exists $Mu^{\top}=c$ and $\|u\|=O_{\delta,k}(1)$. Note that $Mu^{\top}=c$ is equivalent to
@@
To prove Theorem 4.1, it suffices to show the following lemma, as incorporated in Subsection B.4 of the appendix.